Solve dy/dx +2xy= 1+x²+y² given y'(x)=x is a solution.
1
Expert's answer
2020-10-05T11:39:11-0400

"\displaystyle\begin{aligned}
\frac{\mathrm{d}y}{\mathrm{d}x} + 2xy &= 1 + x^2 + y^2 \\
\frac{\mathrm{d}y}{\mathrm{d}x} &=1 + x^2 + y^2 - 2xy\\
&= 1 + (x - y)^2
\end{aligned}\\

\therefore \frac{\mathrm{d}y}{\mathrm{d}x} = 1 + (x - y)^2 \\

\textsf{Making the substitution} \\
u = x - y \\

\frac{\mathrm{d}u}{\mathrm{d}x} = 1 - \frac{\mathrm{d}y}{\mathrm{d}x}\\


\therefore 1 - \frac{\mathrm{d}y}{\mathrm{d}x} = -(x - y)^2 \Rightarrow \frac{\mathrm{d}u}{\mathrm{d}x} = -u^2\\

\textsf{Seperating the variables, we have;}\\

-\frac{\mathrm{d}u}{u^2} = \mathrm{d}x\\


\textsf{Integrating both sides, we have;}\\

-\displaystyle\int\frac{\mathrm{d}u}{u^2} = \displaystyle\int \mathrm{d}x\\

\frac{1}{u} = x + C\\ 


u = \frac{1}{x + C}\\


x - y = \frac{1}{x + C}\\


(x - y)(x + C) = 1\\


y = x - \frac{1}{x + C}\\


\therefore y = x - \frac{1}{x + C} \hspace{0.1cm} \textsf{is also a solution to the}\\\textsf{first order linear differential equation}\\

\textsf{Given that}\hspace{0.1cm} y = x\hspace{0.1cm}\textsf{is also a solution}\\\textsf{to the D.E, then}\\

y = \left(x - \frac{1}{x + C}\right)+ x \hspace{0.1cm} \textsf{is a solution to the}\\\textsf{first order linear differential equation}\\


\therefore y = 2x - \frac{1}{x + C} \textsf{is a solution to the D.E}."